# AlphaApollo: Orchestrating Foundation Models and Professional Tools into a Self-Evolving System for Deep Agentic Reasoning

### v1

@@ Abstract
We present AlphaApollo, a self-evolving agentic reasoning system that aims to address two bottlenecks in foundation model (FM) reasoning—limited model-intrinsic capacity and unreliable test-time iteration. AlphaApollo orchestrates multiple models with professional tools to enable deliberate, verifiable reasoning. It couples (i) a computation tool (Python with numerical and symbolic libraries) and (ii) a retrieval tool (task-relevant external information) to execute exact calculations and ground decisions. The system further supports multi-round, multi-model solution evolution via a shared state map that records candidates, executable checks, and feedback for iterative refinement. In evaluations on AIME 2024/2025 across multiple models, AlphaApollo delivers consistent gains: $5.15\%$ Average@32 and $23.34\%$ Pass@32 for Qwen2.5-14B-Instruct, and $8.91\%$ Average@32 with $26.67\%$ Pass@32 for Llama-3.3-70B-Instruct. Tool-use analysis shows that more than $80\%$ of tool calls are successfully executed, with consistent outperformance of non-tool baselines, thereby lifting the capability ceiling of FMs. More empirical results and implementation details will be updated at https://github.com/tmlr-group/AlphaApollo.  [^1] This project is ongoing. We welcome feedback from the community and will frequently update our technical report.
@@ 1 Introduction
(a) The Apollo Program (in 1960s) for moon landing with humans
(b) The AlphaApollo System (ours) for problem solving with foundation models
Figure 1: The Apollo Program mobilized more than 400,000 people over the span of a decade, from Apollo 1 to Apollo 17, to develop the mission systems that enabled humanity’s first moon landings. As a tribute to this historic achievement, we name our project AlphaApollo. The two projects share common principles despite their different domains: (1) reliance on advanced tools—the Apollo mission system and, in our case, Python code with retrieval systems; (2) collaboration among many participants—people or models; and (3) iterations across a series of missions or solutions.
@@
Foundation models (FMs) increasingly boost diverse applications through explicit reasoning, decomposing complex tasks into steps for more reliable decisions (Li et al., 2025). Yet even state-of-the-art models struggle on challenging benchmarks: as of October 2025, GPT-5 and Gemini 2.5 Pro reach only $25.3\%$ and $21.6\%$ on Humanity’s Last Exam (Phan et al., 2025), and $9.9\%$ and $4.9\%$ on ARC-AGI-2 (Chollet et al., 2025), with open-source models lagging further. Beyond math and code, limited computational power and insufficient domain knowledge constrain their impact in domains such as biology, chemistry, and healthcare, undermining reliability for real-world scientific use.
The two key bottlenecks in FM reasoning are (i) model-intrinsic capacity and (ii) test-time iteration. First, prompting and post-training methods ultimately depend on the base model’s capacity, raising the question of whether observed gains reflect genuinely emergent reasoning or improved elicitation. Behaviors such as self-reflection (“aha” moments) largely originate in pre-training and can be further elicited, whereas abilities requiring exact calculus or symbolic manipulation remain constrained by next-token prediction (Yang et al., 2024c; Wang et al., 2025). Second, effective test-time reasoning often traverses a large solution space with trustworthy feedback. Yet, current scaling strategies—parallel, sequential, or hybrid—heavily rely on the model’s own signals, which are subjective and unreliable (Gao et al., 2025). In the absence of ground-truth verification, reliably assessing solution quality at inference time remains a significant challenge. Moreover, extending iteration to multiple interacting models incurs additional compute and coordination overhead while often yielding little or no benefit over non-interactive single-model scaling.
@@
This work introduces AlphaApollo, an agentic reasoning system designed to overcome these bottlenecks. Its design principle is to orchestrate diverse models and professional tools into a self-evolving system, enabling parallel, deep, agentic reasoning. Inspired by the Apollo program of the 1960s (Figure 1(a)), AlphaApollo adopts a strategy of setting clear goals, concentrating expertise and resources, and coordinating systematic collaboration under shared trust and organizational support, making it possible to tackle complex problems once deemed impossible. To transcend the capacity limits of a single model, AlphaApollo centers on two core features, detailed as follows.
  • First, AlphaApollo couples two classes of professional tools—(i) computation and (ii) retrieval—to enable deliberate, tool-augmented reasoning. The computation tool is a Python interpreter with domain libraries (e.g., SciPy (Virtanen et al., 2020) and SymPy (Meurer et al., 2017)) for numerical and symbolic calculations. The retrieval tool surfaces task-relevant information from external sources (e.g., library documentation or search engines), such as the usage of a specific SciPy function. This agentic paradigm departs from single-model and multi-model paradigms by integrating exact computation with targeted retrieval to strengthen FM reasoning (Figure 2).
  • Second, AlphaApollo enables multi-round, multi-model solution evolution (Figure 3). Each model has full access to the toolset to propose candidate solutions and perform follow-up evaluation. For example, solving a math problem may produce a textual derivation with Python code; the code is then executed and tested to yield verifiable, fine-grained feedback that guides further refinement. All candidates and evaluations are recorded as states in an evolving map (as a shared memory), over which all models operate in parallel, referencing prior states to generate refined solutions.
Figure 2: A comparison of three reasoning paradigms.
Figure 3: Test-time evolving with multiple models.
Empirically, we evaluate AlphaApollo on mathematics reasoning benchmarks, AIME 2024 and 2025, across multiple model families, including Qwen2.5 (7B/14B/32B/72B), Qwen3-235B-A22B, and Llama-3.3-70B-Instruct. Across all models, AlphaApollo yields consistent gains—up to $9.16\%$ Average@32@32 and $23.34\%$ Pass@32@32 on medium Qwen2.5 models; substantial improvements on Llama3.3-70B-Instruct for AIME 2025 ($16.67\%$ Average@32@32; Pass@32@32 increasing from $23.33\%$ to $46.67\%$); and stable gains even on the strong Qwen3-235B-A22B—demonstrating effective scalability. Tool-use analysis further shows that tool-call correctness is near or above $80\%$ for most settings and that tool-augmented responses consistently outperform non-tool ones, indicating that AlphaApollo both improves average performance and raises the capability ceiling for complex reasoning.
AlphaApollo is under active development. This technical report will be updated as new features arrive. The current release introduces our first feature: tool-augmented reasoning. Next, we will add the second feature of test-time scaling—evolving solutions with single or multiple models—followed by broader integration of frontier models, professional tools, and advanced algorithms. We will open-source the full code and experimental results to enable reproducibility and extension.
@@ 2 AlphaApollo
This section presents the technical design of AlphaApollo. Section 2.1 introduces the rollout framework, followed by the computation and retrieval modules in Sections 2.2 and 2.3, respectively.
Tool-augmented reasoning.
As illustrated in Figure 4, AlphaApollo structures the reasoning pipeline (of a single model) into three types of tokens: think, tool call, and tool response. Specifically, the think tokens represent the model’s internal reasoning process. Whenever the model requires external support—computational or informational—it issues a tool call, which is intercepted and executed by the AlphaApollo system. The resulting tool response is then inserted back into the model’s context, allowing the model’s reasoning to resume. Through iterative cycles of thinking, tool calls, and tool responses, the reasoning proceeds until a final solution (with an answer) is generated.
2.1 The Rollout Framework
Figure 4: Schematic illustrating the tool-calling process of the rollout framework. The example shown demonstrates computational tool calls, while the dashed path indicates retrieval tool calls.
In AlphaApollo, the reasoning trajectory generated by a particular model encompasses both model inference (within the think and tool call tokens) and tool execution (within the tool response tokens). This tool-augmented reasoning trajectory is generated by the rollout framework, which bridges the gap between a foundation model and multiple functional tools. Specifically:
@@ 2.1 The Rollout Framework
  • For model inference, AlphaApollo can generate the next token locally or remotely, with a particular model. The supported inference backends include SGLang (Zheng et al., 2024), vLLM (Kwon et al., 2023), HuggingFace Transformers (Huggingface, 2025), and external APIs (OpenAI, 2025).
  • For tool execution, AlphaApollo adopts the Model Context Protocol (MCP) (Anthropic, 2024), which standardizes how tools provide context to foundation models. Through MCP, AlphaApollo seamlessly integrates diverse tools and foundation models under a unified protocol.
Following the MCP’s architecture, AlphaApollo’s rollout framework consists of one manager, two clients, and two servers, as illustrated in Figure 4. Here, each computational or retrieval tool is paired with a server and a client. Specifically:
  • Functions. The manager coordinates the overall rollout, assigning a tool call to a client. The client monitors tool-call status and results, communicating directly with the server. The server executes the tool call and delivers the result to the client, which returns the tool response to the model.
  • Running-time inference. When the model generates a tool call enclosed within <tool_call> and </tool_call> tags, the manager captures it. The manager then assigns the call to the appropriate client, which forwards it to the designated server. The server executes the requested task and returns the result to the client. The client extracts execution results, parses potential errors, and delivers them back to the model, wrapped within <tool_response> and </tool_response> tags.
  • Initialization. The manager loads two configuration files: (1) a server configuration file specifying connection details such as URLs and authentication, and (2) a tool configuration file defining tool-level parameters such as rate limits, timeouts, and tool names. The manager uses the server configuration to launch servers, and the tool configuration to initialize clients. In addition, the model should know all the available tools before reasoning. This is achieved by the tool schemas that define the accessible tools and their usage (details in Appendix A.2), including the tool names, descriptions, and input parameters, together with the system prompt through the chat template.
AlphaApollo guides the model to decide when to invoke a computational or retrieval tool through prompting: the system provides descriptions of available tools and requests the model to determine when and how to call them. Future versions will optimize it via post-training for adaptive tool use.
2.2 The Computational Module
Figure 5: The pipeline of the computational module in processing code with or without errors. The light-blue components (Code checking, Result checking) correspond to Client 1, while the purple components correspond to the Computational Server.
  • SymPy (Meurer et al., 2017) is a computer algebra system for symbolic mathematics, enabling the manipulation and solution of mathematical expressions in closed form. For example, it can determine the exact roots of the cubic equation $x^{3}-2x+1$ within Python.
  • NumPy (Harris et al., 2020) is a fundamental library for fast, vectorized numerical computing, providing powerful support for $N$-dimensional arrays and matrices. Typical tasks include linear algebra operations such as computing the dot product of two matrices.
  • SciPy (Virtanen et al., 2020) is a comprehensive library of advanced numerical algorithms for science and engineering, supporting integration, ordinary differential equations, optimization, signal processing, sparse linear algebra, statistics, and more. For instance, it can readily solve problems like finding the minimum of $f(x)=\sin(x)+x^{2}$ over the interval [3,3][-3,3].
Python environment and code execution.
Error correction.
The model-generated code can contain errors. To improve the robustness against these errors, AlphaApollo’s computational module incorporates a hybrid error-correction mechanism that combines rule-based and model-based approaches. When the model generates Python code, the rule-based approach detects and automatically corrects errors identifiable through predefined rules, and then passes the corrected code to the Python interpreter for execution. Then, if a run-time error arises, the model-based approach analyzes the execution results to identify potential causes and propose targeted correction strategies, thereby guiding the model toward improved subsequent code generation. We outline the specific error types addressed by each approach (more in Appendix B.1):
  • Rule-based error correction. All IndentationError can be detected, with most being automatically correctable. The rule-based approach verifies the legality of indentation line by line, identifying and removing unnecessary space tokens to ensure proper indentation throughout the code. Additionally, certain SyntaxError, such as extraneous markdown blocks wrapping the code, can be detected and resolved. This method extracts the core Python code from wrapped content, ensuring the generated code is executable.
  • Model-based error correction. When run-time errors such as NameError or IndexError occur, this approach provides detailed feedback derived from the execution results, including likely causes and suggested fixes, to guide the model in self-correction. For instance, in the case of a NameError, the feedback may include guidance: “This error is caused by using a module or a variable that is not defined, please check if you have imported all using Python code packages in your code, and you write complete code in your current tool calling.”
In addition, for errors such as ValueError, TypeError, AttributeError, and ImportError, which commonly arise when using external libraries (e.g., calling sympy.factorial with a negative argument, although it only accepts non-negative values), the computational module can invoke the retrieval module (Section 2.3) for further guidance. By combining this capability with the two-fold error-correction mechanism, the computational module creates a model-friendly Python environment that empowers foundation models to perform code-augmented reasoning.
2.3 The Retrieval Module
  • Query rewriter. The workflow begins with the query rewriter, which transforms the initial query into a retrieval-friendly specification. Instead of passing detailed task descriptions directly, the rewriter abstracts them into generalizable forms that emphasize functional intent. For example, an over-detailed query “Solve the equation $12x^{2}-xy-6y^{2}=0$ for $x$ symbolically” is rewritten as a more suitable one, “Solve a quadratic equation symbolically for a variable.” This reformulation abstracts away numerical details while retaining the core intent, resulting in more relevant retrieval terms for the retriever. We implement the query rewriter using an instruction-following model with a tailored prompt template. Details are provided in Appendix A.3.
  • Document retriever. After rewriting the query, the document retriever searches for relevant information in the module’s indexed corpus, which includes the source code of a Python library and its associated documentation. Technically, we partition the corpus into overlapped chunks, which could improve retrieval effectiveness by aligning units with sentence-level semantics and reducing the inclusion of irrelevant content. We implement overlapped chunking with a fixed-length sliding window, allowing adjacent chunks to share overlapping tokens and thus preserve cross-boundary context. These chunks are then encoded by the sentence-level embedding model and stored in a vector database. When the retriever receives a refined query from the rewriter, it first encodes the query using the same embedding model employed during document indexing. The retriever then conducts a cosine-similarity search over the database to locate the top-$K$ chunk embeddings relevant to the query. The corresponding document segments are retrieved and assembled into a composite context, which is then passed to the summarization module.
  • Result summarizer. Finally, the result summarizer filters and summarizes the retrieved context into a concise response. Its role is to highlight callable functions, required arguments, and minimal working examples rather than returning raw documentation. For instance, when the retriever provides (i) demonstrations of sympy.solve, (ii) module descriptions of solveset, and (iii) general API notes on sympy, the summarizer distills them into an actionable tool description. It identifies solve as the appropriate function, specifies its key arguments, and generates a compact example such as: from sympy import solve, symbols; a, b, c, x = symbols(‘a b c x’); solve(ax**2 + bx + c, x). The summarizer is implemented using an instruction-following model with a specialized prompt template, which is provided in Appendix A.3.

### v2

# AlphaApollo: A System for Deep Agentic Reasoning

@@ Abstract
We present AlphaApollo, an agentic reasoning system that targets two bottlenecks in foundation-model reasoning: (1) limited reasoning capacity for complex, long-horizon problem solving and (2) unreliable test-time evolution without trustworthy verification. AlphaApollo orchestrates models and tools via three components: (i) multi-turn agentic reasoning, which formalizes model-environment interaction with structured tool calls and responses; (ii) multi-turn agentic learning, which applies turn-level reinforcement learning to optimize tool-use reasoning while decoupling actions from tool responses for stable training; and (iii) multi-round agentic evolution, which refines solutions through a propose–judge–update loop with tool-assisted verifications and long-horizon memory. Across seven math reasoning benchmarks and multiple model scales, AlphaApollo improves performance through reliable tool use ($85\%$ tool-call success), substantial gains from multi-turn RL (Avg@32: Qwen2.5-1.5B-Instruct $1.07\%\!\rightarrow\!9.64\%$, Qwen2.5-7B-Instruct $8.77\%\!\rightarrow\!20.35\%$), and improvements from evolution (e.g., Qwen2.5-3B-Instruct $5.27\%\!\rightarrow\!7.70\%$, Qwen2.5-14B-Instruct $16.53\%\!\rightarrow\!21.08\%$). This project is still ongoing.  [^1] Project website: https://alphaapollo.org/. We welcome feedback from the community and will frequently update the source code and technical report.
@@ 1 Introduction
Foundation models (FMs) increasingly power diverse applications via explicit reasoning, decomposing complex tasks into manageable steps. Yet single-model reasoning remains insufficient for frontier problems or real-world tasks. As of December 2025, GPT-5 achieves $25.3\%$ and Gemini 2.5 Pro $21.6\%$ on Humanity’s Last Exam (phan2025humanity), and $9.9\%$ and $4.9\%$ on ARC-AGI-2 (chollet2025arc). Beyond math and code, limited compute and domain knowledge restrict performance in biology, chemistry, and healthcare, undermining reliability for scientific discovery.
Two bottlenecks limit FM reasoning are (i) model-intrinsic capacity to generate candidate solutions and (ii) test-time evolution to refine them. First, prompting and post-training are often used to enhance model capacity, but they largely depend on the base model’s priors, making it unclear whether observed gains reflect emergent capabilities or mere elicitation from pre-training (e.g., step-by-step reasoning and self-reflection). Core reasoning skills (e.g., exact calculus and symbolic manipulation) remain constrained by next-token prediction (yang2024number; wang2025hierarchical). Second, without ground-truth verification, test-time evolution often relies on the model’s own judgments, which can be subjective and unreliable (gao2025survey). Finally, scalable parallel evolution remains under-explored, limiting efficiency and multi-model coordination, while long-horizon evolution cannot be achieved without an effective mechanism of memory.
@@
This work introduces AlphaApollo, an agentic reasoning system designed to overcome these bottlenecks. Its design principle is to orchestrate models and tools into a self-evolving system for deep agentic reasoning (as in Fig. 1). AlphaApollo adopts a strategy of setting clear goals, concentrating expertise and resources, and coordinating systematic collaboration under organizational support, thereby enabling it to tackle complex problems. Specifically, AlphaApollo centers on three features to transcend the capacity limits of a single model:
  • Multi-turn agentic reasoning. AlphaApollo solves tasks through iterative model-environment interaction: the model produces a structured action (tool call or answer), the environment executes tools and returns feedback, and the accumulated history serves as dynamic memory.
  • Multi-turn agentic learning. AlphaApollo post-trains at the turn level, optimizing the model’s actions (reasoning and tool invocation) while decoupling them from tool responses, which stabilizes RL/SFT and improves tool-use decisions (what to call, what to query, when to stop).
  • Multi-round agentic evolution. At test time, AlphaApollo iteratively refined solutions with a propose-judge-update loop, using tool-assisted checking and selective memory to refine candidates over multiple rounds. This evolution can coordinate multiple models in solving one problem.
Figure 1: An overview of the AlphaApollo System for problem-solving with foundation models.
Empirically, we evaluate AlphaApollo on seven mathematical reasoning benchmarks across multiple model scales and settings. AlphaApollo achieves consistent gains from agentic reasoning alone (e.g., for Avg@32, Qwen2.5-3B: $4.66\%\!\rightarrow\!4.72\%$; Qwen2.5-14B: $10.82\%\!\rightarrow\!13.49\%$), driven by reliable tool use with over $85\%$ tool-call success across datasets. Multi-turn RL further boosts Avg@32 substantially (Qwen2.5-1.5B: $1.07\%\!\rightarrow\!9.64\%$; Qwen2.5-3B: $4.72\%\!\rightarrow\!13.35\%$; Qwen2.5-7B: $8.77\%\!\rightarrow\!20.35\%$). Finally, test-time agentic evolution delivers additional scalable gains (Qwen2.5-3B: $4.79\%\!\rightarrow\!6.92\%$; Qwen2.5-7B: $9.24\%\!\rightarrow\!10.79\%$; Qwen2.5-14B: $16.53\%\!\rightarrow\!21.08\%$), demonstrating reliable tool use, effective learning, and iterative self-improvement.
@@ 2 AlphaApollo
Figure 2: AlphaApollo’s rollout is an iterative model-environment interaction. The model outputs an action, and the environment provides feedback. Each turn’s trajectory is used as the prompt for the next turn, enabling dynamic memory. The final answer is produced when no tool calls are made.
2.1 Multi-turn Agentic Reasoning
AlphaApollo structures the agentic reasoning as a multi-turn interaction between model and environment (Figure 2). In each turn, the model $\pi_{\theta}$ performs reasoning and invokes a tool call if needed, while the environment captures the tool call, executes the corresponding tool, and provides the tool responses back to the model. Specifically, a trajectory $\tau_{t}$ at $t$-th interacttion turn incorporates prompt $p_{t}$ (the model’s input), model output $o_{t}$, and environment feedback $f_{t}$, namely, $\tau_{t}=(p_{t},o_{t},f_{t})$:
  • The prompt $p_{t}$ incorporates essential information from previous t1t-1 turns. For example, in a three-turn interaction, we have $p_{1}$ (the initial prompt), $p_{2}=(p_{1},o_{1},f_{1})$, and $p_{3}=(p_{2},o_{2},f_{2})$.
  • The output $o_{t}$ consists of the thinking process, followed by either a tool call or the final answer.
  • The feedback $f_{t}$ contains the tool response after executing the tool call.
This model-environment interaction terminates in two situations: (1) the model produces a final answer or (2) reaches the predefined maximum number of turns $T_{\max}$. We denote the number of interacted turns in a reasoning trajectory before termination as $T$ ($T\leq T_{\max}$), and the final-turn trajectory as $\tau_{T}$. Details of the model-environment synergies are introduced as follows.
Environment side.
The environment is responsible for (1) hosting tools, (2) parsing the output $o_{t}$, (3) executing the tool call in $o_{t}$, and (4) returning the feedback $f_{t}$ to the model. Specifically:
  • Hosting tools. Currently, AlphaApollo provides two types of tools: computational tools to solve precise calculations and retrieval tools to retrieve necessary knowledge for problem solving (detailed implementations are introduced in Appendix B). These tools are implemented as callable functions for the model, while users can easily plug in custom tools for extension.
  • Parsing outputs. The environment parses the model’s output $o_{t}$ to identify two contents: (1) tool calls wrapped within tool-specific tags (e.g., <python_code> and </python_code> for computational tools, <local_rag> and </local_rag> for retrieval tools) and (2) the final answer wrapped within <answer> and </answer> tags. The interaction terminates if the final answer is extracted.
  • Executing tools. The system routes each tool call (parsed in the above step) to the corresponding tool hosting in the system. In AlphaApollo, each trajectory is handled by a separate environment, so tool calls across different trajectories are executed asynchronously, enabling efficient parallelism.
  • Returning feedback. The results of tool execution are wrapped within <tool_response> and </tool_response> tags and returned to the model as the feedback $f_{t}$.
Model side.
Given prompt $p_{t}$, the model $\pi_{\theta}$ generates output $o_{t}$ and receives environment feedback $f_{t}$. The system updates the memory and constructs the next-turn prompt $p_{t+1}$. Specifically:
@@ 2.1 Multi-turn Agentic Reasoning
  • Inference infrastructure. AlphaApollo can generate the next token locally or remotely, with a particular model. The supported inference backends include vLLM (kwon2023efficient), SGLang (zheng2024sglang), HuggingFace Transformers (Huggingface), and external APIs (openaiapi). AlphaApollo employs Ray (moritz2018ray) to parallelize trajectory generation. For a batch of questions, the system spawns multiple environments, each handling the interaction for one trajectory, and all environments run concurrently until all trajectories are collected.
  • Output generation. The generated output $o_{t}$ follows a structured format: it begins with a reasoning trajectory where the model articulates its thought process (enclosed within <think> and </think> tags), followed by either a tool call or the final answer. This structured format aligns with the parsing patterns on the environment side, ensuring seamless model-environment interaction.
  • Memory management. The prompt in each turn includes three types of information: the task description that specifies the user query and output format, tool specifications that describe the available tools and their usage formats, and interaction history that records previous outputs and feedback from earlier turns if available. Notably, AlphaApollo supports flexible memory strategies: by default, it concatenates all previous interactions as context, while for long-horizon tasks, it supports long-term memory that selectively retains high-quality trajectories (details in Sec. 2.3).
2.2 Multi-turn Agentic Learning
Figure 3: Illustration of multi-turn agentic RL in AlphaApollo. During generation, per-turn rewards are assigned based on model outputs and environment feedback, and summed to form the trajectory reward. During policy update, the policy is updated at each turn with non-model outputs masked.
Based on the turn-level agentic reasoning system (Sec. 2.1) that decouples model-generated content from tool responses, AlphaApollo incorporates VeRL (sheng2024hybridflow) (an open-source framework for RL) into a turn-level optimization for stable and flexible agentic learning, along with versatile SFT/RL/parameter-efficient algorithms for versatile learning configurations.
Turn-level optimization.
The full trajectory $\tau_{T}$ comprises multiple turns of model-generated content and the environment feedback. Notably, optimizing on the environment feedback can introduce instability (jin2025search; feng2025group). We discuss the trajectory-level optimization in Appendix C. For stable learning on the multi-turn interaction, AlphaApollo adopts turn-level optimization. This paradigm decouples model generations $o_{t}$ from environment feedback $f_{t}$, thereby facilitating targeted optimization on $o_{t}$. For example, the turn-level GRPO is formulated as:
$\mathcal{J}_{\text{Turn-GRPO}}(\theta)\;=\;\mathbb{E}_{(q,a)\sim\mathcal{D},\{o^{(i)}\}_{i=1}^{N}\sim\pi_{\text{old}}(\cdot|q)}\Big[\frac{1}{N}\sum_{i=1}^{N}\frac{1}{T^{(i)}}\sum_{t=1}^{T^{(i)}}\frac{1}{|o^{(i)}_{t}|}\sum_{k=1}^{|o^{(i)}_{t}|}\min\Big(\frac{\pi_{\theta}\big(o_{t,k}^{(i)}\mid p_{t},o_{t,<k}^{(i)}\big)}{\pi_{\text{old}}\big(o_{t,k}^{(i)}\mid p_{t},o_{t,<k}^{(i)}\big)}\,A_{t,k}^{(i)},$ (1)
$\operatorname{clip}\big(\frac{\pi_{\theta}\big(o_{t,k}^{(i)}\mid p_{t},o_{t,<k}^{(i)}\big)}{\pi_{\text{old}}\big(o_{t,k}^{(i)}\mid p_{t},o_{t,<k}^{(i)}\big)},1-\epsilon,1+\epsilon\big)\,A_{t,k}^{(i)}\Big)-\beta\,\mathbb{D}_{\mathrm{KL}}\bigl[\pi_{\theta}\,\big\|\,\pi_{\mathrm{ref}}\bigr]\,\Big],$
where $o_{t,k}^{(i)}$ denotes $k$-th token of the model’s output in $t$-th turn of the $i$-th trajectory, $N$ is the number of group samples for GRPO, $A_{t,k}^{(i)}=\frac{r(\tau_{T^{(i)}})-\bar{r}}{\sigma_{r}}$ where $r$ is the reward by evaluating the answer proposed in turn $T$, $\bar{r}$ and $\sigma_{r}$ are the baseline and the normalization across all $N$ trajectories of one question. $A$ indicates the advantage and $\epsilon$ is the clipping bound, and $\beta\,\mathbb{D}_{\mathrm{KL}}[\pi_{\theta}\,\big\|\,\pi_{\mathrm{ref}}]$ is the KL divergence term using an unbiased estimator (kl_approx) weighted by $\beta$. Here, token importance ratios $\nicefrac{{\pi_{\theta}(o_{t,k}^{(i)}\mid p_{t},o_{t,<k}^{(i)})}}{{\pi_{\text{old}}(o_{t,k}^{(i)}\mid p_{t},o_{t,<k}^{(i)})}}$ are computed conditioned on $p_{t}$ and generated tokens $o_{t,<k}$, which incorporate the query and prior turns’ generations and tool responses.
Similarly, the turn-level SFT is formulated as:
$\mathcal{J}_{\text{Turn-SFT}}(\theta)=\mathbb{E}_{\tau_{T}\sim\mathcal{D}_{\text{SFT}}}\left[\sum_{t=1}^{T}\sum_{k=1}^{|o_{t}|}\log\pi_{\theta}(o_{t,k}\mid p_{t},o_{t,<k})\right],$ (2)
where $\tau_{T}$ is a full trajectory curated from the multi-turn agentic reasoning (Sec. 2.1).
Post-training features.
To create a versatile and user-friendly agentic learning system, AlphaApollo incorporates the following post-training features:
  • Algorithms: AlphaApollo supports common SFT/RL algorithms for post-training, such as PPO (schulman2017proximal), GRPO (shao2024deepseekmath), and DAPO (yu2025dapo).
  • Models: AlphaApollo enables agentic learning across various foundation model families, including Qwen2.5 (yang2024qwen2), Qwen3 (yang2025qwen3), and Llama3.2 (grattafiori2024llama). Additionally, it provides a multi-modal preprocessing and training pipeline for post-training with multi-modal foundation models, such as Qwen2.5-VL families (bai2025qwen2).
  • Flexible advantage estimation: AlphaApollo supports both trajectory- and turn-level advantage estimation in RL, making it compatible with emerging process reward modeling algorithms.
  • Parameter-efficient fine-turning: Leveraging FSDP and vLLM as backends, AlphaApollo enables LoRA (hu2022lora) for efficient post-training of large-scale models.
2.3 Multi-round Agentic Evolution
Figure 4: Illustration of multi-round agentic evolution in AlphaApollo. The model iteratively refines its strategies through a propose-judge-update evolutionary loop. A long-term memory is introduced to prevent future errors and promote efficient strategies in subsequent rounds.
AlphaApollo employs a test-time evolution mechanism to iteratively refine solutions (Figure 4). Operating via a propose-judge-update loop, the system coordinates multiple agents to generate solutions, evaluate them, and store the judgment in long-term memory to guide subsequent rounds.
Overview.
We denote the system’s initial input as $p_{1}$ and the long-term memory as $\mathcal{M}$, initialized as $\mathcal{M}^{(1)}=\emptyset$. The process iterates over multiple rounds $r=\{1,\dots,R\}$. In each round, the system coordinates four specialized agents $\pi$ to instantiate the pipeline, yielding a candidate solution $\hat{s}^{(r)}$ and diagnostic judgment $j^{(r)}$. These outputs are formed as a memory entity $(\hat{s}^{(r)},j^{(r)})$, which is committed to $\mathcal{M}$. The $\mathcal{M}^{(r+1)}$ is retrieved from $\mathcal{M}$ to condition the policy $\pi$ for the subsequent evolution. Detailed prompts and implementations can be found in Appendix D.2. The pipeline is formulated as follows:
1. Propose: $\quad\tau^{(r)}\sim\pi_{\text{solver}}(\cdot\mid p_{1},\mathcal{M}^{(r)}),\quad\hat{s}^{(r)}\sim\pi_{\text{abs}}(\tau^{(r)});$ (3)
2. Reflection: $\quad\mathcal{V}^{(r)}=\{v_{i}\}_{i=1}^{N}\sim\pi_{\text{eval}}(\cdot\mid\hat{s}^{(r)}),\quad j^{(r)}\sim\pi_{\text{sum}}(\mathcal{V}^{(r)});$
3. Update: $\quad\mathcal{M}=\mathcal{M}\cup\{(\hat{s}^{(r)},j^{(r)})\},\quad\mathcal{M}^{(r+1)}=\mathrm{Retrieve}(\mathcal{M}).$
We detail the role of each agent in this pipeline below.
  • Solver ($\pi_{\text{solver}}$): Given a problem and memory state, generates a multi-turn reasoning trajectory. The solver interacts with the environment to generate a complete trajectory $\tau^{(r)}$.
  • Abstractor ($\pi_{\text{abs}}$): Given $\tau^{(r)}$, generates a condensed solution. The abstractor compresses $\tau^{(r)}$ into a condensed solution $\hat{s}^{(r)}$ to fit the context window while preserving essential information.
  • Evaluator ($\pi_{\text{eval}}$): Given $\hat{s}^{(r)}$, generates verifications $\mathcal{V}^{(r)}$. For deterministic tasks, it integrates execution feedback. For open-ended tasks, it aggregates $N$ sampled critiques via majority voting.
  • Summarizer ($\pi_{\text{sum}}$): Given $\mathcal{V}^{(r)}$, generates comprehensive judgement $j^{(r)}$. Synthesize $\mathcal{V}^{(r)}$ into a high-level judgment $j^{(r)}$, remove redundancy to provide clean advice for $\pi_{\text{solver}}$ in future rounds.
Long-term memory for evolution.
To enable long-horizon evolution, the memory module supports efficient retrieval to prevent excessive context, including the overlong content of correctness and solution. Specifically, the long-term memory operates in two phases:
  • Store: At the end of round $r$, the system transitions the memory state by appending the new entity: $\mathcal{M}\cup\{(\hat{s}^{(r)},j^{(r)})\}$. This accumulation ensures the history of exploration is preserved.
  • Retrieval: To yield $\mathcal{M}^{(r+1)}$, the system selects the Top-$K$ entries from the memory under a weighted scoring function prioritizing (1) the correct solution and (2) shorter solutions in correct groups.
Retrieving high-level judgment $j$ promotes promising strategies while preventing recurrent errors. Furthermore, the module employs thread-safe locks to decouple storage from execution, enabling asynchronous agents to reliably synchronize insights regardless of latency.
Parallel Evolution.
To accelerate reasoning, AlphaApollo distributes the evolution pipeline across multiple worker threads. This architecture supports heterogeneous solvers (e.g., mixing open-source models with proprietary APIs or varying sampling temperatures). Agents synchronize via the shared long-term memory: when one solver commits a high-quality solution, it becomes available to others, fostering collective intelligence that guides the evolution toward a better solution.